# Performance Comparison of HTTP/3 and HTTP/2: Proxy vs. Non-Proxy Environments

### v1

@@ 4.2. Multi-Stream Experiment Evaluation
The resulting graph in Figure 5 shows the performance of H2 and H3 under various loss levels with 20 streams. The data indicates that when the packet loss is below 8%, H2 and H3 exhibit comparable performance. However, when the packet loss reaches or exceeds 8%, H3 significantly outperforms H2, with the performance gap widening as the loss percentage increases. This demonstrates H3’s superior resilience to high packet loss conditions.
@@ 4.3. Connection Migration Experiment Evaluation
Figure 6 displays the results of the connection migration experiments, comparing the performance of H3 and H2 clients under various network conditions. These results highlight the advantages of QUIC’s connection migration feature. The H3 client (quiche_h3) consistently exhibited lower total request times across all scenarios involving IP address changes.
Across all tested network conditions, H2 clients lacking the TCP range request capability (e.g., wget_h2_no_resume and curl_h2_no_resume) showed significantly higher total request times and were more adversely affected by IP changes and poor network conditions. However, when the TCP range request feature was employed (wget_h2_with_resume), H2’s performance became comparable to H3 in scenarios involving either packet loss or delay alone. While the TCP range request feature can be advantageous for specific use cases like file downloads, H3 is generally more suited for a broader range of applications, including real-time environments and the noisy network conditions often encountered in mobile scenarios.
@@
Analysis The primary advantage of H3’s connection migration feature is rooted in QUIC’s ability to seamlessly manage changes in network paths, such as those that occur when a device transitions between Wi-Fi and cellular networks. QUIC’s connection migration allows an ongoing session to continue without interruption, preserving the state of the connection even when the client’s IP address changes. This capability extends to proxy-enhanced environments, where H3 continues to maintain stable performance despite the additional complexity introduced by the proxy. In contrast, TCP-based H2 requires a new connection to be established following an IP change, leading to higher latency and increased total request times. This issue is especially noticeable in clients that do not support the TCP range request feature.
The results indicate that H3’s connection migration not only maintains the session but also minimizes total request times, which is essential for real-time applications where reducing delays is critical. This efficiency is a direct result of QUIC’s design, which includes mechanisms to quickly re-establish connection parameters and resume data transmission following an IP change. The fact that H3 rarely experiences failed migrations, even under high-loss and high-delay conditions, further underscores its robustness.
@@
$4.$ Web Page Loading Experiment Evaluation
@@ 4.5. General Observations
3. Role of Congestion Control: The experiments also underscore the impact of CCAs on protocol performance. While both H2 and H3 can benefit from advanced CCAs like BBR, H3’s internal congestion control mechanisms appear more adept at handling varying network conditions. The use of BBR in proxies significantly boosts H2’s performance by reducing latency and optimizing throughput. However, H3’s inherent design makes it less reliant on external congestion management.
@@ 5. Future Work
Future research should focus on exploring the interaction of different CCAs with HTTP/3 across various network environments. Beyond BBR, testing algorithms like Cubic, NewReno, and newer ones could reveal how they affect QUIC’s performance, offering insights into optimizing configurations for specific use cases such as real-time applications, video streaming, and mobile networks. Examining these factors could improve the efficiency and adaptability of H3 in diverse scenarios.
Additionally, investigating the role of proxies in optimizing web protocols, especially in edge computing and CDN environments, would be valuable. Future work could enhance H3 support in proxies and develop specific optimizations for QUIC, particularly for high-latency or lossy networks. Moreover, studies on security and privacy implications related to QUIC’s fast connection establishment and migration could assess their impact on data integrity and confidentiality in scenarios like man-in-the-middle attacks or network handoffs. This research is essential for securing H3 in sensitive environments and large-scale deployments under real-world traffic conditions.
@@ A.1. Clients and Server Protocols and Implementations
The clients and server protocols and implementations show in Table 2.

### v2

@@ 4.2. Multi-Stream Experiment Evaluation
The resulting graph in Figure 5 shows the performance of H2 and H3 under various loss levels with 20 streams. The data indicates that when the packet loss is below 8%, H2 and H3 exhibit comparable performance. However, when the packet loss reaches or exceeds 8%, H3 significantly outperforms H2, with the performance gap widening as the loss percentage increases. This demonstrates H3’s superior resilience to high packet loss conditions. Specifically, at 12% packet loss without a proxy, H2 required an average of 113.42 seconds to complete the download, whereas H3 completed the same task in just 20.98 seconds, representing an 81.5% improvement.
@@ 4.3. Connection Migration Experiment Evaluation
Figure 6 displays the results of the connection migration experiments, comparing the performance of H3 and H2 clients under various network conditions. These results highlight the advantages of QUIC’s connection migration feature. The H3 client (quiche_h3) consistently exhibited lower total request time across all scenarios involving IP address changes. Notably, under conditions of 4% packet loss and 50ms delay, the H2 client (wget_h2_resume) took an average of 56.7 seconds to complete the download following an IP address change without a proxy, whereas the H3 client (quiche_h3) required only 6.6 seconds. This represents an 88.36% improvement in performance.
Across all tested network conditions, H2 clients lacking the TCP range requests capability (e.g., wget_h2_no_resume and curl_h2_no_resume) showed significantly higher total request time and were more adversely affected by IP changes and poor network conditions. However, when the TCP range requests feature was employed (wget_h2_resume), H2’s performance became comparable to H3 in scenarios involving either packet loss or delay alone. While the TCP range requests feature can be advantageous for specific use cases like file downloads, H3 is generally more suited for a broader range of applications, including real-time environments and the noisy network conditions often encountered in mobile scenarios.
@@
Analysis The primary advantage of H3’s connection migration feature is rooted in QUIC’s ability to seamlessly manage changes in network paths, such as those that occur when a device transitions between Wi-Fi and cellular networks. QUIC’s connection migration allows an ongoing session to continue without interruption, preserving the state of the connection even when the client’s IP address changes. This capability extends to proxy-enhanced environments, where H3 continues to maintain stable performance despite the additional complexity introduced by the proxy. In contrast, TCP-based H2 requires a new connection to be established following an IP change, leading to higher latency and increased total request time. This issue is especially noticeable in clients that do not support the TCP range requests feature.
The results indicate that H3’s connection migration not only maintains the session but also minimizes total request time, which is essential for real-time applications where reducing delays is critical. This efficiency is a direct result of QUIC’s design, which includes mechanisms to quickly re-establish connection parameters and resume data transmission following an IP change. The fact that H3 rarely experiences failed migrations, even under high-loss and high-delay conditions, further underscores its robustness.
@@
[…]
Web Page Loading Results.
4.4. Web Page Loading Experiment Evaluation
@@ 4.5. Discussion
3. Role of Congestion Control: The experiments also underscore the impact of CCAs on protocol performance. While both H2 and H3 can benefit from advanced CCAs like BBR, QUIC’s internal congestion control mechanisms appear more adept at handling varying network conditions. The use of BBR in proxies significantly boosts H2’s performance by reducing latency and optimizing throughput. However, H3’s inherent design makes it less reliant on external congestion management.
@@ 5. Future Work
Future research should focus on exploring the interaction of different CCAs with QUIC across various network environments. Beyond BBR, testing algorithms like Cubic, NewReno, and newer ones could reveal how they affect QUIC’s performance, offering insights into optimizing configurations for specific use cases such as real-time applications, video streaming, and mobile networks. Examining these factors could improve the efficiency and adaptability of H3 in diverse scenarios.
Additionally, investigating the role of proxies in optimizing web protocols, especially in edge computing and CDN environments, would be valuable. Future work could enhance H3 support in proxies and develop specific optimizations for H3, particularly for high-latency or lossy networks. Moreover, studies on security and privacy implications related to QUIC’s fast connection establishment and migration could assess their impact on data integrity and confidentiality in scenarios like man-in-the-middle attacks or network handoffs. This research is essential for securing H3 in sensitive environments and large-scale deployments under real-world traffic conditions.
@@ A.1. Clients and Server Protocols and Implementations
The clients and server protocols and implementations are shown in Table 2.